OpenStudy (anonymous):

2x + y = 3 -3x + 4y = 1 [6.02] Use the Substitution Method to solve the following system of equations.

OpenStudy (tkhunny):

Well, go ahead and show us your best work.

OpenStudy (anonymous):

Alright well you have a fairly easy value in order to obtain your substitution; the first value set isolate y; once y is isolated substitute into the second equation.

OpenStudy (anonymous):

y=3-2x -3x+4)3-2x)=1 -3x+12-8x=1 -11x+12=1 -11x+12-12=1=1-12 -11x=-11 -11x/11=-11/-11 x=1

OpenStudy (anonymous):

2x+y=3 2(1)=y=3 2+y=3 2-2+y=3-2 y=1 (1,1) is that correct?

OpenStudy (texaschic101):

2x + y = 3 ---> y = 3 - 2x -3x + 4y = 1 -3x + 4(3 - 2x) = 1 -3x + 12 - 8x = 1 -11x + 12 = 1 -11x = - 11 x = 1 2x + y = 3 2(1) + y = 3 2 + y = 3 y = 3 - 2 y = 1 You are correct....the answer is (1,1) By the way, you can check your answer by subbing in known values into either one, or both, of the equations and if they come out equal, you are correct. Let me show you... -3x + 4y = 1 2x + y = 3 -3(1) + 4(1) = 1 OR 2(1) + 1 = 3 -3 + 4 = 1 2 + 1 = 3 1 = 1 (correct) 3 = 3 (correct)

OpenStudy (anonymous):

Thanks!

OpenStudy (texaschic101):

anytime :)